A man with a mass of 90 kg is skiing at a speed of 20 m.s-1 when he reaches the 

bottom of a 20 m slope. He goes up the slope without exerting any other forces to 

help him. Ignore friction. Will he reach the top of the slope? Show all calculations.


1
Expert's answer
2021-10-15T10:23:57-0400

The law of conservation of energy gives

"\frac{mv_0^2}{2}=mgh"

So

"h=\frac{v_0^2}{2g}=\frac{20^2}{2*9.8}=20.4\:\rm m"

Since "h>20 \;\rm m", so the man will reach the top of the slope.